OpenStudy (anonymous):

a/7+1/4=6 linear equations

OpenStudy (anonymous):

\[\frac{a}{7}+\frac{1}{4}=6\]?

OpenStudy (anonymous):

yes thats it

OpenStudy (anonymous):

subtract \(\frac{1}{4}\) from both sides, then multiply both sides by \(7\)

OpenStudy (anonymous):

in other words, first find \(6-\frac{1}{4}\) once you have that number, multiply it by \(7\)

OpenStudy (anonymous):

10.5

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

i really doubt that what is \[6-\frac{1}{4}\]?

OpenStudy (anonymous):

1.5

OpenStudy (anonymous):

you have six dollars you spend one quarter you do not have a dollar fifty left

OpenStudy (anonymous):

if you have $6 and spend $0.25 how much is left?

OpenStudy (anonymous):

5.75

OpenStudy (anonymous):

thats better !

OpenStudy (anonymous):

multiply that by \(7\) and you are done

OpenStudy (anonymous):

but can you how me how you got 5.75?

OpenStudy (anonymous):

you said it, not me

OpenStudy (anonymous):

its 40.25 but really can you tell me

OpenStudy (anonymous):

btw the final answer is not \(5.75\) that is just \(6-\frac{1}{4}\) the final answer is \[7\times 5.75\]

OpenStudy (anonymous):

look at the thread above i never said it was 5.75, you did i only agreed so however you did it must be correct

OpenStudy (anonymous):

you didnt show me the step to get that you said if you have 6$ and spend 0.25 how much is left.

OpenStudy (anonymous):

the first piece of arithmetic was six minus one quarter right?

OpenStudy (anonymous):

it is pretty obvious that six minus one quarter is five and three quarters

OpenStudy (anonymous):

ohhh ok i get it